OpenStudy (anonymous):

Consider the functions f(x) = 2x + 1 and g(x) = x2 − 10. What is the value of f[g(3)]? 39 −3 −7 -1

OpenStudy (nincompoop):

you start from the inside then work your way through outside

OpenStudy (nincompoop):

first, solve g(x) since you are being asked to solve g(3) in the innermost function.

OpenStudy (anonymous):

x^3-10?

OpenStudy (anonymous):

-10?

OpenStudy (anonymous):

@pooja195

Nnesha (nnesha):

g(3) here x =3 so replace x by 3 not the exponent...

OpenStudy (anonymous):

so (3)-10?

Nnesha (nnesha):

and where is exponent 2 ??

OpenStudy (anonymous):

(3)^2-10?

Nnesha (nnesha):

\[\huge\rm g(\color{reD}{x}) = \color{red}{x}^2 -10\]

Nnesha (nnesha):

yes right!

Nnesha (nnesha):

and then substitute g(3) into f(x) function

OpenStudy (anonymous):

yikes, idk how

Nnesha (nnesha):

so what is g(3) =??

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @lealealea222 (3)^2-10? \(\color{blue}{\text{End of Quote}}\) =what ?

OpenStudy (anonymous):

-1

OpenStudy (nincompoop):

so what is g(3) ?

OpenStudy (anonymous):

i think -1

OpenStudy (nincompoop):

okay now use f(x) being that x is -1

OpenStudy (anonymous):

is the answer -3 ?

OpenStudy (nincompoop):

look at this and tell me what you observe given: f(x) = 2x + 1 and g(x) = x2 − 10 \(f(x) = 2(x^2-10) +1 \)

OpenStudy (nincompoop):

tick tock tick tock lol

OpenStudy (anonymous):

man...

OpenStudy (anonymous):

idk haha

OpenStudy (nincompoop):

did you notice that I put the g(x) function into the f(x) function

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

look at it again

OpenStudy (anonymous):

you put them together

OpenStudy (anonymous):

i think its -3

OpenStudy (nincompoop):

how so what is \(2 \times (-1) \)

OpenStudy (anonymous):

-2x

OpenStudy (nincompoop):



OpenStudy (nincompoop):

it's not variable x, but the multiplication operation

OpenStudy (nincompoop):

use the number line I drew and place where -2 is

OpenStudy (anonymous):

oh ok let me do that again

OpenStudy (anonymous):

-2

OpenStudy (nincompoop):

use the number line I drew and place where -2 is

OpenStudy (nincompoop):

click the pencil icon on the upper right hand corner of my drawing so you can use it

OpenStudy (anonymous):

 ?